# P2LHAP:Wearable sensor-based human activity recognition, segmentation and forecast through Patch-to-Label Seq2Seq Transformer

### v1

@@ I Introduction
The intersection of deep learning and wearable sensor technology for real-time human activity recognition has emerged as a prominent research focus in recent years. This area holds significant promise for applications in medical health, assisted living, smart homes, and other domains. In various practical scenarios, there is a need for perceptual methods to delineate activity boundaries, identify activity categories, and predict forthcoming activities from the continuous stream of sensor data. As a result, human activity perception encompasses three pivotal tasks: human activity segmentation (HAS)[51], activity recognition (HAR)[50], and activity forecast (HAF)[49]. While extensive research[42, 43] has addressed each of these tasks individually, limited attention has been devoted to simultaneously addressing all three tasks in a cohesive framework.
@@
Several approaches exist for concurrently achieving activity segmentation, identification, and forecast, as illustrated in Fig. 1. The first approach Fig. 1(a) employs three independent single-task learning models in parallel. These models output the category of the current activity, the boundary of the current activity, and the category of the next activity. However, this method has drawbacks, including disparate data annotation requirements for different models, the need for additional policy compatibility model outputs, and a lack of information sharing between models, leading to inefficient use of computing resources.
@@
A second approach involves multi-task learning in Fig. 1(b), where multiple interrelated tasks share data and encoders, learning a common data representation. Subsequently, multiple lightweight single-task models [5, 6, 25] are employed. While this approach often enhances multitasking effectiveness, it introduces increased model complexity compared to single-task models. Managing this complexity may necessitate additional computational resources and careful architectural design. Furthermore, these multitask models can be more sensitive to hyperparameter choices, demanding meticulous tuning to achieve optimal performance across all tasks.
@@
Different from the first two methods, as shown in Fig. 1(c), this paper proposes a novel Patch-to-Label Seq2Seq framework that tackles all the three human activity perception tasks in a single-task model, named P2LHAP. P2LHAP adopts the encoder-decoder framework, the input is a patch sequence of sensor data, and the output is a sequence of patch-level activity label, where a patch is a short segment of sensor data points. With patch-level labels, as shown in Fig. 1(c), it can simply Merge adjacent identical labels to determine the category of the active boundary. In addition, using the encoder-decoder architecture of seq2seq, It is easy to output future activity token sequences, making it simple to predict activity label. The architecture handling capability of variable-length sensor activity sequences is excellent, and it exhibits high adaptability to sequences of indefinite length. Moreover, it comprehensively considers all information within the input sensor signal sequence, effectively retaining the contextual information of the sensor data stream.
@@
Using patches offers several advantages over individual data points and sliding windows. Compared to using single data points[16, 6], using patches can reduce the impact of noise, allowing feature extraction within local intervals[41]. Most existing activity recognition research uses one or more fixed-length slides Windows segment the data stream, assuming each window contains a complete activity[5]. However, due to the changing duration of the activity, the window length is very difficult to set. If it is too small, the window does not contain enough information; if it is too large, a window contains activities of multiple different categories, both of which will reduce the recognition accuracy [53]. The patching strategy can circumvent this problem well, because we assume that an activity can contain any number of patches, and encoder-decoder architectures such as tranfomer are good at handling variable-length sequences, making it feasible to identify activities of arbitrary length.
@@
We introduce a channel-independent transfomer as the encoder-decoder of P2LHAP. Although transformers have achieved great results in the fields of NLP[7] and CV[55], recent research shows that classic transformers do not work well on multi-variable long sequence time series prediction problems[54]. This may be because multi-variable features will interfere Extraction of long-term sequence features[41]. Human activity perception based on wearable devices is essentially a long-sequence multi-variable time series analysis problem. Sensor data includes acceleration, gyroscopes, magnetometers, etc. channel signal. Inspired by[41], instead of using the classic transformer architecture, we introduce a channel-independent transformer architecture as the encoder-decoder of P2LHAP, each patch only contains data from a single wearable sensor channel information, each sensor signal sequences within the dataset undergoes processing independently within the Transformer model, allowing for the learning of diverse attention patterns tailored to each sequence. In contrast to the conventional architecture where all sequences are subject to the same attention mechanism, this individualized approach prevents potential interference caused by disparate behaviors within multi-channel sensor data. Moreover, by adopting a channel-independent strategy, the adverse impact of noise prevalent in certain sensor signal sequences can be mitigated as it no longer gets projected onto other sequences within the embedding space.
we utilize the original sensor data sequence along with the patch features acquired by the encoder to compute the representation of individual patches using the attention network for the prediction of activity labels assigned to each patch. Simultaneously, due to the variable size of the patch, the ultimate active label sequence may suffer from over-segmentation. To counter this issue, we devised a smoothing algorithm for the final Patch-level active label sequence, aimed at mitigating the over-segmentation phenomenon. This algorithm considers the number of activity categories surrounding the patch to adjust the current patch-level activity label.
@@
  • The concept of transforming patch sequences into label sequences was introduced for conducting activity perception tasks, where activity recognition, segmentation, and forecast are performed on the resulting activity label sequence. This approach has demonstrated enhanced performance across all three tasks.
  • Introduced a channel-independent Transformer architecture that can transform the Patch level sensor sequence at the current moment into an active label sequence corresponding to both the current and future Patch blocks.
@@
  • The proposed P2LHAP method has been evaluated through experiments conducted on three widely used benchmark datasets for activity recognition, segmentation, and forecast. The experimental results demonstrate that our approach surpasses existing methods in terms of performance.
@@ II-A Sensor-based activity recognition and segmentation tasks
Sensor-based human activity recognition (HAR) research primarily relies on inertial measurement unit (IMU) sensors found in wearable devices such as smart watches, smart insoles, and sports straps. The common approach for sensor-based activity recognition involves utilizing fixed-size sliding windows, which has shown impressive results[8]. [45] a long short-term memory (LSTM)-based classification method is proposed to identify fine-grained patterns according to the high-level features contextually extracted from the sequential motion data. Alternatively, [32] formulates HAR as a set forecast problem, where the prediction consists of ongoing active elements with varying cardinality. In the realm of HAR, [33] introduces an attention-based multi-head model for human activity recognition. Likewise, [46] propose a semi-supervised generative adversarial network (GAN) for CSI-based activity recognition (CsiGAN). Furthermore, [35] presents a triplet cross-dimensional attention method for sensor-based activity recognition tasks, deploying three attention branches to capture cross-interactions among sensor dimensions, time dimensions, and channel dimensions. Similarly, [36] introduces a novel multi-branch CNN that incorporates an attention mechanism to perform kernel selection among multiple branches with different receptive fields in the HAR domain. Additionally, [10] suggests a GAN-based approach to synthesize sensor data for sensor-based HAR tasks. In the context of realistic settings, [11] and [12] utilize sample evaluations to compare the performance of their recognition models. However, the length of the sliding window directly impacts the quality of feature extraction and classification accuracy [8]. Consequently, researchers have started focusing on adaptive sliding windows and have proposed sensor data segmentation strategies based on sliding windows. For instance, [13] introduces a novel multi-window fusion method that involves segmenting data using five different window sizes and fusing the recognition results in two stages. In a different approach, [14] presents a unified weakly supervised framework based on distributed kernel embeddings. This framework segments sensor streams, extracts powerful features from each segment, and trains a final classifier for activity recognition. Additionally, [15] suggests using a multivariate Gaussian distribution to detect signal differences between time windows. Nonetheless, defining the optimal window size remains a challenge due to the variation in activity durations.
Another effective method for sensor-based activity recognition is dense label prediction [4]. Inspired by the image semantic segmentation task, [18, 19, 16, 17] apply FCN (fully convolutional network) and U-net to sensor-based activity recognition to predict dense labels, achieving good recognition effects. [20] further proposes Conditional-UNet for coherent human activity recognition with multiple dense labels, explicitly modeling the conditional dependencies between these labels. Lastly, [33] explores the exploitation of potential relationships among multi-channel sensor modes and specific activities, investigating the effectiveness of data-agnostic enhancement of multi-modal sensor data streams for activity recognition prediction.
@@
Sensor-based human activity recognition systems have gained popularity with the widespread use of wearable devices in daily life. However, there is a notable lack of research focusing on sensor-based activity forecast. Recent studies have introduced attention models to predict multivariate motion signals from IMU sensors, thereby preserving pattern and feature information across multiple channels. Predictive models for these time series data can be utilized for human activity forecast by generating and classifying future activity signals. In [21], a combination of a neural network and Fourier transform was employed to convert the sensor activity data into an image format for predicting future signals. The system achieved a mean square error (MSE) value of 0.134 and was successfully used to predict elderly falls. In [22] introduced the SynSigGAN model, which utilizes bidirectional grid long short-term memory as the generator network of the GAN model, and a convolutional neural network as the discriminator network to generate future signals. Furthermore, recent work [23] proposes a Transformer-based GAN that can successfully generate real synthetic time series data sequences of arbitrary length, similar to real data sequences. Both the generator and discriminator networks of the GAN model are built using a pure Transformer encoder architecture, which outperforms traditional GAN structures based on RNN or CNN models. An additional study [41] presented the PatchTST time series forecast model, which is based on channel independence and patching, and demonstrated promising results. Although progress has been made in multivariate sensor time series data forecast, research on multi-category forecast of human activities based on multi-sensor signal channels remains limited.
@@ II-C Sensor-based multi-task learning framework
Sensor-based multi-task activity recognition presents a significant challenge in the field. [24] employs recurrent neural networks to segment and identify activities and cycles using inertial sensor data. [25] proposes an efficient multi-task learning framework that utilizes commercial smart insoles. This framework addresses three health management-related problems by converting smart insole sensor data into recursive graphs and employing an improved MobileNetV2 as the backbone network. Experimental results demonstrate that this multi-task learning framework outperforms single-task models in tasks such as activity classification, speed estimation, and weight estimation. [26] introduces a multi-task model based on LSTM for predicting each activity and estimating activity intensity using wearable sensor signals. [6] combines the segmentation and classification tasks by passing the sensor signal through the MS-TCN module to predict sample-level labels and activity boundaries. This joint segmentation and classification approach calculates sample-level loss and boundary consistency loss to achieve improved performance. [5] proposes a novel multi-task framework called MTHARS, which leverages the dynamic characteristics of activity length to segment activities of varying durations through multi-scale window splicing. This approach effectively combines activity segmentation and recognition, leading to improved performance in both tasks.
@@ III Method
In this section, we introduce our proposed method, which is a multi-task model for human activity recognition utilizing the Patch-to-Label Seq2Seq approach. Fig. 2 provides an overview of the model’s architecture. This model enables activity recognition, segmentation, and forecast tasks on continuous sensor data streams collected from wearable devices. Our approach consists of three primary components: 1) Patch-to-Label Seq2Seq framework, 2) Patching and channel independent Transformer architecture, and 3) activity classification, segmentation, and forecast module.
@@ III-A Patch-to-Label Seq2Seq framework
[…]
@@ III-B1 Patching and channel independent method
We process each sensor signal channel independently, specifically, We segment the input sensor signal sequence $\mathbf{X}_{1:L}$ into $M$ channel sensor signal sequences $\mathbf{x}^{(i)}$ based on its dimension $M$, $\mathbf{x}^{(i)}$ represents the sensor sequence of the $i$-th channel, $i\in\{1,2,\cdots,M\}$. we normalize each $\mathbf{x}^{(i)}$ before patching and the mean and deviation. Next, we divide $\mathbf{x}^{(i)}$ into overlapping or non-overlapping Patches and the length of the Patch block is $P$. The overlapping part is its step size, represented by $S$. By segmentation, a patch-level sensor sequence $\mathbf{x}_{p}^{(i)}\in\mathbb{R}^{1\times N}$ will be obtained, where $N=\left\lfloor\frac{(L-P)}{S}\right\rfloor+2$. $\mathbf{x}_{p}^{(N)}$ is the last patch of the segmented patch sequence, this patch is learned through the previous $N-1$ patches, a sequence representation into the future is obtained, with its initial value filled by repeatedly replicating the last number of the original sequence. Finally, we add the corresponding position coding to $\mathbf{x}_{p}^{(i)}$ to obtain $\mathbf{x}_{d}^{(i)}$, and then transmit the time series of each channel to the Transformer backbone respectively.

### v3

@@ I Introduction
The intersection of deep learning and wearable sensor technology for real-time human activity recognition has emerged as a prominent research focus in recent years. This area holds significant promise for applications in medical health, assisted living, smart homes, and other domains. In various practical scenarios, there is a need for perceptual techniques to delineate activity boundaries, identify activity categories, and predict forthcoming activities from the continuous stream of sensor data. As a result, the field of human activity perception encompasses three pivotal tasks: human activity segmentation (HAS)[1], activity recognition (HAR)[2], and activity forecast (HAF)[3]. While extensive research[4, 5] has addressed each of these tasks individually, limited attention has been devoted to simultaneously addressing all three tasks in a cohesive framework.
@@
Several approaches exist for concurrently achieving activity segmentation, recognition, and forecast, as illustrated in Fig. 1. The first approach, shown in Fig. 1(a), employs three independent single-task learning models in parallel. These models output the category of the current activity, the boundary of the current activity, and the category of the next activity. However, this method has drawbacks, including disparate data annotation requirements for different models, the need for additional policies to ensure compatibility among model outputs, and a lack of information sharing between models, leading to inefficient use of computing resources.
@@
A second approach involves multi-task learning in Fig. 1(b), recent advancements in multi-task Human Activity Recognition learning frameworks [6, 7, 8] leverage the interdependencies among tasks to facilitate data sharing and encoder learning, enabling the extraction of common data representations. Following this, multiple lightweight single-task models are employed. While this approach often enhances multitasking effectiveness, it introduces increased model complexity compared to single-task models. Managing this complexity may require additional computational resources and careful architectural design. Simultaneously, it is crucial to ensure consistency in the model output representation across different tasks. Our objective is to standardize the output representation for improved task-specific adaptability. Furthermore, these multitask models can be more sensitive to hyperparameter choices, requiring meticulous tuning to achieve optimal performance across all tasks.
@@
Different from the first two methods, as shown in Fig. 1(c), this paper proposes a novel Patch-to-Label Seq2Seq framework that tackles all the three human activity perception tasks in a single-task model, named P2LHAP. P2LHAP adopts the encoder-decoder framework, the input is a patch sequence of sensor data, and the output is a sequence of patch-level activity label, where a patch is a short segment of sensor data points. With patch-level labels, it can simply Merge adjacent identical labels to determine the category of the active boundary. In addition, using the encoder-decoder architecture of seq2seq, it can easily output future activity token sequences, making it simple to predict activity label. The architecture handling capability of variable-length sensor activity sequences is excellent, and it exhibits high adaptability to sequences of indefinite length, comprehensively considering all information within the input sensor signal sequence and effectively retaining the contextual information of the sensor data stream.
@@
Using patches offers several advantages over individual data points and sliding windows. Compared to using single data points[10, 7], patch-based methods help mitigate the effects of noise by enabling feature extraction within localized intervals[11]. Most existing activity recognition research employs one or more fixed-length sliding Windows to segment the data stream, under the assumption that each window captures a complete activity[6]. However, because activities have varying durations, it is very difficult to set an appropriate window length. If it is too short, the window may not capture sufficient information; if it is too large, may encompass multiple activities of different categories, both of which will reduce the recognition accuracy [9]. The patching strategy can circumvent this problem well, because we assume that an activity can contain any number of patches, and encoder-decoder architectures such as Tranfomer are good at handling variable-length sequences, making it feasible to identify activities of arbitrary duration.
@@
Although Transformers have achieved great results in the fields of NLP[13] and CV[14], recent research shows that classic Transformers do not work well on multi-variable long sequence time series prediction problems[15]. This may be because multi-variable features will interfere Extraction of long-term sequence features[11]. Limited training data is also a issue, in contrast to image and text data, the scarcity of activity recognition datasets can impact the Transformer model’s performance. Despite their modeling flexibility surpassing that of convolutional neural networks (CNNs) or recurrent neural networks (RNNs), Transformers can pinpoint relevant information at any position. Consequently, substantial datasets are typically necessary to realize their full potential. Human activity perception based on wearable devices is essentially a long-sequence multi-variable time series analysis problem. Sensor data includes signals from accelerometers, gyroscopes, magnetometers, etc. Inspired by[11], instead of using the classic Transformer architecture, we introduce a channel-independent Transformer architecture as the encoder-decoder of P2LHAP, each patch only contains data from a single wearable sensor channel information, each sensor signal sequences within the dataset undergoes processing independently within the Transformer model, allowing for the learning of diverse attention patterns tailored to each sequence. In contrast to the conventional architecture where all sequences are subject to the same attention mechanism, this individualized approach prevents potential interference caused by disparate behaviors within multi-channel sensor data. Moreover, by adopting a channel-independent strategy, the adverse impact of noise prevalent in certain sensor signal sequences can be mitigated as it no longer gets projected onto other sequences within the embedding space.
we utilize the original sensor data sequence along with the patch features obtained from the encoder to compute patch-level representations through an attention network, which are then used to predict activity labels for each patch. However, due to the variable size of the patch, the ultimate active label sequence may suffer from over-segmentation. To address this issue, we devised a smoothing algorithm for the final Patch-level activity label sequence, aimed at mitigating the over-segmentation phenomenon. This algorithm considers the number of activity categories surrounding the patch to adjust the current patch-level activity label. It then establishes activity boundaries based on the label range, significantly reducing boundary ambiguity in the activity signal.
@@
  • We proposes a novel Patch-to-Label Seq2Seq framework that tackles all the three human activity perception tasks in a single-task model, named P2LHAP. To the best of our knowledge, our work constitutes the pioneering effort in unifying activity segmentation, recognition, and forecast within a single framework.
  • We employ patches to address the challenge of insufficient information from large sliding windows or individual data points, and our method, utilizing a seq2seq encoder-decoder architecture, readily generates future activity token sequences, making it straightforward to forecast activity labels.
  • Each patch of our approach only contains data from a single wearable sensor channel information, each sensor signal sequences within the dataset undergoes processing independently within the Transformer model, allowing for the learning of diverse attention patterns tailored to each sequence.
@@
  • The proposed P2LHAP method has been evaluated through experiments conducted on three widely used benchmark datasets for activity recognition, segmentation and forecast. The experimental results demonstrate that our approach surpasses existing methods in terms of performance.
@@ II-A Sensor-based activity recognition and segmentation tasks
Sensor-based human activity recognition primarily relies on IMU sensors in wearable devices like smartwatches, smart insoles, and exercise bands. Traditional approaches, such as fixed-size sliding windows and dense labeling, have demonstrated strong performance [16]. Recent studies have employed diverse techniques. [17] distinguished static and dynamic activities using statistical techniques, then classified specific activities with random forests (RF) and CNNs. In[18], NAS was used to investigate effective deep learning (DL) architectures for smartphone inference. [19] introduced a lightweight CNN-BiGRU model, leveraging inertial sensor data from smartwatches and smartphones, achieving satisfactory results. However, the inherent uncertainty in activity duration, these methods will lead to multi-class window problems, where multiple activities may coexist within a window, introducing noise and negatively impacting model recognition performance. [20] employed an LSTM-based method to identify fine-grained patterns using high-level features from sequential motion data, [21] introduced a deformable convolutional network for activity recognition from complex sensory data, and [22] presented a GRU-INC model that initializes attention-based GRUs to effectively leverage spatiotemporal information in time series. Additionally, [23] proposed a multi-branch CNN-BiLSTM network that extracts features from raw sensor data with minimal preprocessing. [24] further developed Conditional-UNet, which models conditional dependencies between dense labels for coherent HAR. Current advancements in activity segmentation and recognition, such as [1, 7], leverage dense labeling methods. However, these methods rely solely on individual sensor samples as neural network inputs, and will obtain different representations depending on the task. Our proposed approach employing a patch block, situated between the scales of a large sliding window and a single sample point, for data segmentation. We then utilize a Seq2Seq architecture to unify activity segmentation and recognition tasks, obtaining a unified activity label sequence representation to perform different tasks while mitigating the challenges associated with multi-class windows and over segmentation.
@@
Sensor-based human activity recognition systems have gained popularity with the widespread use of wearable devices in daily life. However, there remains a significant gap in research focusing on sensor-based activity forecasting. Recent studies have introduced attention models to predict multivariate motion signals from IMU sensors, thereby preserving pattern and feature information across multiple channels. Predictive models for these time series data can be utilized for human activity forecast by generating and classifying future activity signals. In [25], a combination of a neural network and Fourier transform was employed to convert the sensor activity data into an image format for predicting future signals. In [26] introduced the SynSigGAN model, which utilizes bidirectional grid long short-term memory as the generator network of the GAN model, and a convolutional neural network as the discriminator network to generate future signals. Furthermore, recent work [27] proposes a Transformer-based GAN that can successfully generate real synthetic time series data sequences of arbitrary length, similar to real data sequences. The GAN model’s generator and discriminator are built using a pure Transformer encoder architecture, which outperforms traditional GAN structures based on RNN or CNN models. Despite these advancements, existing methods predominantly rely on training models to forecast and emulate future sensor activity signals, subsequently employing classification algorithms to derive activity categories from the forecasted signal sequences. Any inaccuracies in the forecasted sensor signals inevitably impact the final classification outcome. In contrast, our proposed P2LHAP architecture bypasses this intermediate step by directly forecasting future activity category token sequences, thereby enhancing forecast accuracy. An additional study [11] presented the PatchTST time series forecast model, which is based on channel independence and patching, and demonstrated promising results. While there has been progress in forecasting multivariate sensor time series data, research on multi-category forecast of human activities based on multi-sensor signal channels remains limited.
@@ II-C Sensor-based multi-task learning framework
Sensor-based multi-task activity recognition presents a significant challenge in the field. [28] employs recurrent neural networks to segment and identify activities and cycles using inertial sensor data. [8] proposes an efficient multi-task learning framework that utilizes commercial smart insoles. This framework addresses three health management-related problems by converting smart insole sensor data into recursive graphs and employing an improved MobileNetV2 as the backbone network. Experimental results demonstrate that this multi-task learning framework outperforms single-task models in tasks such as activity classification, speed estimation, and weight estimation. [29] introduces a multi-task model based on LSTM for predicting each activity and estimating activity intensity using wearable sensor signals. [7] combines the segmentation and classification tasks by passing the sensor signal through the MS-TCN module to predict sample-level labels and activity boundaries. This joint segmentation and classification approach calculates sample-level loss and boundary consistency loss to achieve improved performance. [6] proposes a novel multi-task framework called MTHARS, which leverages the dynamic characteristics of activity length to segment activities of varying durations through multi-scale window splicing. This approach effectively combines activity segmentation and recognition, leading to improved performance in both tasks. However, prior studies employed distinct backbone networks for various tasks or failed to achieve unified task representations, our proposed architecture integrates these tasks within a single model for simultaneous learning and representation. By doing so, we obtain a unified label sequence representation.
@@ III Method
In this section, we introduce our proposed method, which is a multi-task model for human activity recognition utilizing the Patch-to-Label Seq2Seq approach. Fig. 2 provides an overview of the model’s architecture. This model enables activity recognition, segmentation, and forecast tasks on continuous sensor data streams collected from wearable devices. Our approach consists of three primary components: 1) Patch-to-Label Seq2Seq framework, 2) Patching and channel independent Transformer architecture, and 3) Activity classification, segmentation, and forecasting module.
@@ III-A Patch-to-Label Seq2Seq framework
The Patch-to-Label Seq2Seq framework we propose is shown in Fig. 2. We define the $M$-dimensional input wearable sensor sequence as $\mathbf{X}_{1:L}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{L}],$ $\mathbf{x}_{t}\in\mathbb{R}^{1\times M}$, where $\mathbf{x}_{t}$ represents the sensor signals collected at timestamp t. The activity label for each sample is denoted as $y_{1:L}=[y_{1},\cdots,y_{L}],$ $y_{t}\in R^{\textit{C}}$, where C represents the number of activity categories, t is the time step, L is the length of the wearable sensor sequence, and the length of the patch block is defined as P. In this study, we do not employ a fixed-length sliding window to segment the sensor data. The number of patches is denoted as L/P. Our approach utilizes a Transformer encoder and decoder as the underlying framework. To simultaneously process the sensor patch of each channel in parallel within the Transformer backbone, it is crucial to duplicate the Transformer weights multiple times and employ the patching operator to create a sensor sequence sample of a size $B\times M\times L$. The patch-level sequence extends to $M\times P\times N$. Reshaping the dimension sequence into a 3D tensor of size $(B\cdot M)\times P\times N$, the batch of samples serves as the input for the encoder. The purpose of the encoder is to map multiple sequences of single-channel sensor data patches into a latent representation. Following this, the decoder is used to derive the representation of the activity sequence, denoted as $\mathbf{Z}^{{}^{\prime}(i)}$. The ultimate output signifies the active label result obtained through the classification head, the active label sequence and the forecast of future activity label sequence of length $T_{pre}$, denoted as $\hat{y}_{pre}=\left\{\hat{y}_{1},\cdots,\hat{y}_{\textit{L}+T_{pre}}\right\}$.
@@ III-B1 Patching and channel independent method
We process each sensor signal channel independently, specifically, we segment the input sensor signal sequence $\mathbf{X}_{1:L}$ into $M$ channel sensor signal sequences $\mathbf{x}^{(i)}$ based on its dimension $M$, $\mathbf{x}^{(i)}$ represents the sensor sequence of the $i$-th channel, $i\in\{1,2,\cdots,M\}$. We normalize each $\mathbf{x}^{(i)}$ before patching and the mean and deviation. Next, we divide $\mathbf{x}^{(i)}$ into overlapping or non-overlapping Patches and the length of the Patch block is $P$. The overlapping part is its step size, represented by $S$. By segmentation, a patch-level sensor sequence $\mathbf{x}_{p}^{(i)}\in\mathbb{R}^{1\times N}$ is obtained, where $N=\left\lfloor\frac{(L-P)}{S}\right\rfloor+2$. $\mathbf{x}_{p}^{(N)}$ is the last patch of the segmented patch sequence, this patch is learned through the previous $N-1$ patches, a sequence representation into the future is obtained, with its initial value filled by repeatedly replicating the last number of the original sequence. Finally, we add the corresponding position coding to $\mathbf{x}_{p}^{(i)}$ to obtain $\mathbf{x}_{d}^{(i)}$, and then transmit the time series of each channel to the Transformer backbone respectively.